OpenStudy (anonymous):

sum n=1 to infinity of (2/n+7)

OpenStudy (anonymous):

diverges

OpenStudy (anonymous):

meaning \(\infty\)

OpenStudy (anonymous):

We know \(1/n\) diverges. \[ \frac 1n =\frac 2{2n} = \frac 2{n+n} < \frac{2}{n+7} \]